William A. Stein Budget

(617) 308-0144  was@math.harvard.edu  http://modular.fas.harvard.edu

Budget Justification Page

A SENIOR PERSONNEL: The PI is requesting 2/9 of his regular salary. This request is based on current annual salary with 3% increase for each successive year.

B OTHER PERSONNEL: Support for two undergraduates to work on the modular forms database for two months each summer. $6000/summer.

C FRINGE BENEFITS: Fringe benefits for summer salary for the first year will be 22.3%, 23.3% for the second year and 24.3% for the 3rd year. Fringe benefits for undergraduates will be 7.6% for two years and 8.6% for 3rd year.

D EQUIPMENT:

– Web server: Apple Mac G5 dual 2GHz processors with 8GB RAM, 2x250GB Serial-ATA hard disks, ATI Radeon 9800 Pro graphics adapter, 23” Apple Flat Panel display, including hardware service: $9,894.00. This computer will be the web server that supports the modular forms database. It will also be used to extend the tables in the database.

– Disk array: Apple Xserve RAID mass storage device with 2.52 Terabytes disk space (14x180GB Ultra ATA disks), single Apple Fibre Channel PCI host controller, enhanced RAID Controller Cache and Cache Battery Backup, including hardware service: $11,732.00. This is where the modular forms database will be stored.

• TRAVEL:

1. Domestic: Visit UC Berkeley periodically to meet with collaborators (Ken Ribet on textbook, Robert Coleman, and Bjorn Poonen): $1500. Attend the winter AMS meeting: $600/year.

2. Foreign: Every other summer, I usually travel to Germany (Oberwolfach), France, and Spain to attend summer conferences and work at the Max Planck Institute: $3000/visit.

• OTHER DIRECT COSTS:

1. Material and Supplies:

– Money for scientific books which will be required to conduct my research. For example, money to purchase books that describe new computer technology that may be used to enhance the modular forms database. $500/year.

William A. Stein Budget

(617) 308-0144  was@math.harvard.edu  http://modular.fas.harvard.edu

– Portable computer: $2500 for a Thinkpad T40 laptop, which will be used by the PI for interactive presentations of research results, development of Magma software and the web interface to the modular forms database, and complex modular forms computations during presentations or when internet access is not available to the PI.

2. Modular curves seminar: Support for the modular curves seminar, which I organize, to help defer the expenses of outside speakers. $1000/year